# Open the Eyes of MPNN: Vision Enhances MPNN in Link Prediction

### v2

@@ Abstract
Message-passing graph neural networks (MPNNs) and structural features (SFs) are cornerstones for the link prediction task. However, as a common and intuitive mode of understanding, the potential of visual perception has been overlooked in the MPNN community. For the first time, we equip MPNNs with vision structural awareness by proposing an effective framework called Graph Vision Network (GVN), along with a more efficient variant (E-GVN). Extensive empirical results demonstrate that with the proposed frameworks, GVN consistently benefits from the vision enhancement across seven link prediction datasets, including challenging large-scale graphs. Such improvements are compatible with existing state-of-the-art (SOTA) methods and GVNs achieve new SOTA results, thereby underscoring a promising novel direction for link prediction. The official code is available at https://github.com/WEIYanbin1999/EGVN.
@@ 1 Introduction
Link prediction is a core task in graph machine learning that focuses on forecasting the presence of a connection between two nodes. This task is crucial in numerous applications such as recommendation systems (He et al., 2020), drug interaction prediction (Yamanishi et al., 2008), and knowledge-based reasoning (Bordes et al., 2013; Wei et al., 2023).
Message-passing graph neural networks (MPNNs) (Kipf & Welling, 2016; Hamilton et al., 2017) are prominent and powerful tools for link prediction. They generate node representations and aggregate them into link representations to predict the existence of links. In addition to MPNNs, various structural features (SFs), which are characteristics derived from the graph topology to capture the relationships and properties of nodes and edges, are integrated into MPNNs to enhance their link prediction capabilities (Zhang & Chen, 2018a; You et al., 2021; Zhu et al., 2021; Yun et al., 2021; Chamberlain et al., 2023; Wang et al., 2024). This integration led to significant advancements, making these SF-enhanced MPNNs dominate the link prediction task.
Considering human perception of graph data, one of the most crucial means is intuitively interpreting of graph information through visual perception. While the structural awareness in MPNNs is derived from message-passing paths, and SFs are typically based on specific heuristic priors, both approaches have overlooked the potential of utilizing visual modalities to comprehend graph structures. As a fundamental capability of human perception, vision has matured into a thriving field supported by robust research tools, such as readily available visual encoders such as VGG16 (Simonyan & Zisserman, 2014), ResNet50 (He et al., 2016), and ViT (Dosovitskiy et al., 2021). These tools offer powerful capabilities for visual information awareness and extraction. Various research domains, including natural language processing, have successfully integrated visual modalities, leading to significant advancements in areas such as visual question-answering (Antol et al., 2015) and multimodal large language models (Zhang et al., 2024).
@@
  • RQ2: Does incorporating vision awareness into graph structures genuinely enhance link prediction? If so, what are the underlying reasons?
@@
Main Contributions. (i) We propose novel frameworks GVN and E-GVN, that first incorporate the vision modality into MPNN for link prediction. They are fully compatible with current models, ensuring seamless integrations. (ii) We reveal properties during graph visualization tailored to link prediction tasks and provide good practices. (iii) We analyze the underlying benefits of vision awareness in link prediction and provide empirical demonstrations. (iv) We delve into the scalability on large-scale graphs and propose efficient optimization in E-GVN. (v) We demonstrate the efficacy of GVN and E-GVN across seven common link prediction datasets, where vision awareness brings orthogonal enhancements to SOTA methods.
@@ 2 Related Works
Link Predictor. Link predictors can be divided into three classes: 1) Node embedding methods (Perozzi et al., 2014; Tang et al., 2015; Grover & Leskovec, 2016) represent each node as an embedding vector and utilize the embeddings of target nodes to predict links. 2) Link prediction heuristics (Liben-Nowell & Kleinberg, 2003; Barabási & Albert, 1999; Adamic & Adar, 2003; Zhou et al., 2009) create SFs through manual design. 3) MPNN-based link predictors produce node representations via the message-passing mechanism.
Structural Features. Structural Features (SFs) are characteristics derived from the graph topology to capture the relationships and properties of nodes and edges. Common SFs are divided into two types: 1) Common Neighbor-based SFs, including Common Neighbor Count (CN) (Barabási & Albert, 1999), Resource Allocation (RA) (Zhou et al., 2009) and Adamic-Adar (AA) (Adamic & Adar, 2003). 2) Path-based SFs, such as Shortest Path Distance (SPD) (Dijkstra, 1959), Double Radius Node Labeling (DRNL) (Zhang & Chen, 2018b) and Distance Encoding (DE) (Li et al., 2020).
SF-enhanced MPNNs. The expressive power of naive MPNN architectures is limited (Zhang et al., 2021), constrained by the 1-dimensional Weisfeiler-Lehman (1-WL) test (Morris et al., 2019), and they fail to finely perceive substructures like triangles (Chen et al., 2020). To overcome these limitations, more recent studies adopt various structural features to enhance MPNNs. For example, SEAL (Zhang & Chen, 2018a) incorporates SPD as structural features into MPNNs by concatenating the SPD from each node to the target nodes $u$ and $v$ with the node features. Neo-GNN (Yun et al., 2021) and BUDDY (Chamberlain et al., 2023) use the heuristic function to model high-order common neighbor information. NCNC (Wang et al., 2024) directly concatenates the weighted sum of node representations of common neighbors with the Hadamard product of MPNN representations of $u$ and $v$.
Graph Learning with Vision. Several research explore vision potentials in graph learning. Das et al. (2023) find the vision-language models have competitive node classification ability. Wei et al. (2024) shows vision perception benefits language-based graph reasoning tasks. However, none of the previous works realize the potential of vision in enhancing MPNN for link prediction, that is the focus of this paper.
@@ 3 Preliminaries
Notations. A Graph $G=(V,E)$ comprises of a set $V$ of $n$ nodes and a set $E$ of $e$ links. We denote the adjacency matrix of $G$ by $A\in\mathbb{R}^{n\times n}$, where $A_{(u,v)}>0$ if and only if the edge $(u,v)$ exists (i.e., $(u,v)\in E$). We define $N(v):=\{v|v\in V,A_{uv}>0\}$ as the set of neighbors of node $v$, and $N_{k}(v)$ as the set of neighbors of node $v$ within $k$ hops, where a node $u\in N_{k}(v)$ if and only if $u$ and $v$ are reachable within $k$ hops. The node feature matrix $\mathbf{X_{G}}\in\mathbb{R}^{n\times F}$ contains the node features in $G$, where the $v$-th row $\mathbf{x_{v}}$ corresponds to the feature of node $v$. We use $S^{k}_{uv}=(V_{uv}^{k},E_{uv}^{k})$ to denote a $k$-hop subgraph enclosing the link $(u,v)$, where $V_{uv}^{k}$ is the union of $k$-hop neighbors of $u$ and $v$, and $E_{uv}^{k}$ is the union of links that can be reached by a $k$-hop walk originating at $u$ or $v$. $S^{k}_{v}$ is the $k$-hop subgraph enclosing node $v$.
MPNNs for Link Prediction. The MPNN is a commonly used model for link prediction tasks. In MPNN, the message-passing mechanism is employed to iteratively update node representations based on information exchanged between neighboring nodes. Mathematically, this message-passing mechanism can be written as
@@
where $\bm{h}_{v}^{t}$ is the representation of node $v$ after $t$ layers, $U^{t}$, $M^{t}$ and $\mathbf{AGG}$ are the update, message-passing, and aggregation functions of the layer $t$. $\mathbf{Y_{G}}\in\mathbb{R}^{n\times F^{\prime}}$ is the matrix of final node representations by MPNN for graph $G$, whose $v$-th row $\mathbf{y}_{v}$ is the final representation of $v$. Given $\mathbf{Y_{G}}$, link probabilities can then be computed as $p(u,v)=R(\mathbf{y}_{u},\mathbf{y}_{v})$, where $R$ is a learnable readout function.
@@ 4 Methodology
In this section, we present the proposed GVN framework. An illustration of GVN is presented in Fig. 1. From Section 4.1 to Section 4.3, we will delve into the detailed designs inside, along with addressing the research questions raised in the introduction. In Section 4.4, we propose a computationally efficient variant of GVN for large-scale graphs.
@@ 4.1 Scope-decoupled Subgraph Visualization
Link-centered Subgraph Visualization. Beginning with RQ1 (i.e., For the link prediction task, what is the effective approach to be aware of graph structures from vision?), we transform graph structures into visual representations by rendering them as images during pre-processing. Given large-scale graphs often with thousands of nodes and edges, we utilize the $k$-hop subgraph sampling, a common practice in graph learning. This method concentrates on visual awareness via the most relevant local $k$-hop subgraph $S^{k}_{uv}$ around the central link $(u,v)$, while excluding distant structures that may introduce noise. In line with the locality principle [^1] Locality Principle: Correlated information is contained within the near neighbors (Bronstein et al., 2017; Wu et al., 2020). and effective practices of many MPNNs with $k\leq 3$ (Kipf & Welling, 2016; Hamilton et al., 2017; Xu et al., 2018; Velickovic et al., 2017; Chen et al., 2022), we limit $k$, which defines the visualized scope, to a maximum of 3.
To be specific, utilizing existing automated graph visualization tools (Details in Appendix G) such as Graphviz (Gansner & North, 2000), Matplotlib (Tosi, 2009), and Igraph (Gabor Csardi, 2006), a given query link $(u,v)$ and its surrounding $k$-hop subgraph $S^{k}_{uv}$ are mapped one-to-one into an image $I^{k}_{uv}$ by a graph visualizer $\mathrm{GV}$. Formally, this is expressed as $I^{k}_{uv}=\mathrm{GV}(S^{k}_{uv},u,v)$.
@@
1) Keeping Style Consistency. It is crucial that all visualized subgraphs maintain consistency in the graph visualizer $\mathrm{GV}$, such as uniformly using Graphviz with fixed configurations (e.g., graph layout algorithm, node colors, and shapes). This consistency prevents inconsistent image styles from introducing additional learning difficulties and unstable training.
@@
3) Ablating Node Labels. All the node labels are omitted, enabling the model to focus on the graph structure, which benefits the generalizability.
Figure 2: Illustrations of visual graph images with different values of $k$.
Decoupled Vision Scope. While the perception scope expands during message passing in MPNN, we fix the visual perception scope $k$ during visualization. This approach ensures that the scope in vision can be decoupled from which in message passing and offers two key advantages. First, both theoretical analysis and empirical evidence (Zeng et al., 2021) show that a decoupled scope independent from message-passing can improve the expressivity of MPNN by addressing the over-smoothing issue (Li et al., 2018; Kenta Oono, 2020; Shen et al., 2024), and enhances the computational scalability of the models by alleviating the neighbor explosion problem (Chiang et al., 2019; Zeng et al., 2019). Second, as shown in Figure 2, an excessively large visual perception scope can lead to overcrowding of nodes/edges on a fixed-size canvas, resulting in a cluttered appearance that complicates graph interpretation. By using a small $k$, we maintain image clarity independent of message passing.
@@ 4.2 Adaptive Extraction of Visual Structural Features
With the same perceptive scope, VSFs can discriminate links involving isomorphic nodes, which MPNNs or the 1-dimensional Weisfeiler-Lehman test cannot.
@@
Fine-grained Substructure Awareness. The ability to capture substructures (motifs) is another perspective of MPNN expressive power and plays a crucial role in areas like biology, molecular, and social networks (Chen et al., 2020; Kanatsoulis & Ribeiro, 2024; Yan et al., 2024). However, this can be challenging for MPNNs (Chen et al., 2020; Arvind et al., 2020). We demonstrate in the following that VSFs are helpful for MPNNs in capturing substructures. Specifically, we follow the experimental setup in (Chen et al., 2020). This involves generating two types of graph datasets, including Erdős-Rényi graphs and random regular graphs, and compute the counts of two substructures including triangles and 3-stars as ground-truth labels. We use MPNN followed by a 3-layer MLP decoder as the architecture and the VSFs are concatenated after the MPNN-produced representations. The evaluation metric is the normalized mean squared error (i.e., the ratio of the mean squared error over the variance of ground-truth labels). Table 1 shows the best and median (third-best) performance over five runs with different random seeds. As can be seen, incorporating VSFs significantly improves the substructure capturing abilities of MPNNs. More details of this experiment are provided in Appendix E.
@@
Information-rich and Adaptive SF Extraction. While the design of existing structural features rely on heuristics, VSFs encode the whole subgraph, making them more information-rich. Figure 4 illustrates the reproduction ratios of various SFs based on VSFs on ogbl-ddi, which is defined as the percentage (%) of corresponding SFs successfully predicted by feeding the VSFs extracted from the pretrained ResNet50 into a trainable 3-layer MLP. The results demonstrate that VSFs effectively capture most of the information from a wide array of structural features, underscoring their comprehensive nature as a versatile structural feature pool.
@@
The reproduction ratios of various SFs reflect the relative distribution of information within the VSFs. In Table 2, we compare the reproduction ratios of SFs between VSFs obtained from a pre-trained ResNet50 and those after finetuning. The latter is derived from a ResNet50 encoder fine-tuned for link prediction with a 2-layer MLP on the ogbl-ddi and PubMed datasets. Note that ogbl-ddi is a very dense graph, while PubMed is much sparser. As shown in Table 2, after finetuning on ogbl-ddi, the reproduction ratio of path-based SFs (SPD, DRNL and DE) consistently decrease. This aligns with underperformed efficacy shown in path-based link prediction models like SEAL (Zhang & Chen, 2018b) and NBFNet (Zhu et al., 2021) tend to underperform in ogbl-ddi dataset, where most node pairs are reachable within two hops, making path-based structural features like SPD less informative. On the other hand, in the sparser PubMed dataset, where path information is more valuable, their reproduction ratios increase after finetuning. The reproduction ratios of common neighbor-based SFs (CN, RA, AA) consistently increase, as models like (Yun et al., 2021; Chamberlain et al., 2023; Wang et al., 2024) consistently benefit from common neighbor SFs across various datasets. These observations demonstrate that VSFs can intelligently adjust the internal information types to meet specific demands of the different scenarios through finetuning.
@@
Table 2: Reproduction ratios of SFs in VSFs with and without finetuning on ogbl-ddi and PubMed datasets.
@@
Through finetuning, VSFs can adaptively adjust their emphasis on different types of structural information, providing scenario adaptability.
VSFs are not permutation-equivariant, yet they remain effective with MPNNs, as shown in Section 5.1. This is potentially attributed to two reasons: 1) Learnable VSFs capture specific information like structural patterns or motifs (see Remark 4.2). 2) Visualization introduces permutation-based augmentations (e.g., varying layouts), making the decoder less sensitive to node order.

### v3

# Open Your Eyes: Vision Enhances Message Passing Neural Networks in Link Prediction

@@ Abstract
Message-passing graph neural networks (MPNNs) and structural features (SFs) are cornerstones for the link prediction task. However, as a common and intuitive mode of understanding, the potential of visual perception has been overlooked in the MPNN community. For the first time, we equip MPNNs with vision structural awareness by proposing an effective framework called Graph Vision Network (GVN), along with a more efficient variant (E-GVN). Extensive empirical results demonstrate that with the proposed framework, GVN consistently benefits from the vision enhancement across seven link prediction datasets, including challenging large-scale graphs. Such improvements are compatible with existing state-of-the-art (SOTA) methods and GVNs achieve new SOTA results, thereby underscoring a promising novel direction for link prediction. The official code is available at https://github.com/WEIYanbin1999/EGVN.
@@ 1 Introduction
Link prediction, which predicts the presence of a connection between two nodes, is a core task in graph machine learning. It is crucial in numerous applications such as recommendation systems (He et al., 2020), drug interaction prediction (Yamanishi et al., 2008), and knowledge-based reasoning (Bordes et al., 2013; Wei et al., 2023).
Message-passing graph neural network (MPNN) (Kipf & Welling, 2017; Hamilton et al., 2017) is a prominent and powerful tool for link prediction. It generates node representations and aggregates them into link representations to predict the existence of links. Besides, various structural features (SFs), which are characteristics derived from the graph topology to capture relationships and properties of nodes and edges, are integrated into MPNNs to enhance their link prediction capabilities (Zhang & Chen, 2018; You et al., 2021; Zhu et al., 2021; Yun et al., 2021; Chamberlain et al., 2023; Wang et al., 2024). This integration led to significant advancements, making the SF-enhanced MPNNs dominate the link prediction task.
Considering human perception of graph data, one of the most crucial means is intuitively interpreting of graph information through visual perception. While the structural awareness in MPNNs is derived from message-passing paths, and SFs are typically based on specific heuristic priors, both approaches have overlooked the potential of utilizing visual modalities to comprehend graph structures. As a fundamental capability of human perception, vision has matured into a thriving field supported by robust research tools, such as readily available visual encoders such as VGG16 (Simonyan & Zisserman, 2015), ResNet50 (He et al., 2016), and ViT (Dosovitskiy et al., 2021). These tools offer powerful capabilities for visual information awareness and extraction. Various research domains, including natural language processing, have successfully integrated visual modalities, leading to significant advancements in areas such as visual question-answering (Antol et al., 2015) and multimodal large language models (Zhang et al., 2024).
@@
  • RQ2: Does incorporating vision awareness into graph structures enhance link prediction? If so, what are the underlying reasons?
@@
Main Contributions. (i) We propose novel frameworks GVN and E-GVN, that are the first in incorporating the vision modality into MPNN for link prediction. They are fully compatible with current models, ensuring seamless integrations. (ii) We reveal properties during graph visualization tailored to link prediction tasks and provide good practices. (iii) We analyze the underlying benefits of vision awareness in link prediction and provide empirical demonstrations. (iv) We delve into the scalability on large-scale graphs and propose efficient optimization in E-GVN. (v) We demonstrate the efficacies of GVN and E-GVN across seven common link prediction datasets, where vision awareness brings orthogonal enhancements to the SOTA methods.
@@ 2 Related Works
Link Predictor. Link predictors can be divided into three classes: 1) Node embedding methods (Perozzi et al., 2014; Tang et al., 2015; Grover & Leskovec, 2016) represent each node as an embedding vector and utilize the embeddings of target nodes to predict links. 2) Link prediction heuristics (Liben-Nowell & Kleinberg, 2003; Barabási & Albert, 1999; Adamic & Adar, 2003; Zhou et al., 2009) create SFs through manual design. 3) MPNN-based link predictors produce node representations via the message-passing mechanism (Kipf & Welling, 2017).
Structural Features. Structural Features (SFs) are characteristics derived from the graph topology to capture the relationships and properties of nodes and edges. Common SFs are divided into two types: 1) Common Neighbor-based SFs, including the Common Neighbor Count (CN) (Barabási & Albert, 1999), Resource Allocation (RA) (Zhou et al., 2009) and Adamic-Adar (AA) (Adamic & Adar, 2003). 2) Path-based SFs, such as the Shortest Path Distance (SPD) (Dijkstra, 1959), Double Radius Node Labeling (DRNL) (Zhang & Chen, 2018) and Distance Encoding (DE) (Li et al., 2020).
SF-enhanced MPNNs. The expressive power of naive MPNN architectures is limited (Zhang et al., 2021), constrained by the 1-dimensional Weisfeiler-Lehman (1-WL) test (Morris et al., 2019), and they fail to finely perceive substructures like triangles (Chen et al., 2020). To overcome these limitations, more recent studies adopt various structural features to enhance MPNNs. For example, SEAL (Zhang & Chen, 2018) incorporates SPD as structural features into MPNNs by concatenating the SPD from each node to the target nodes $u$ and $v$ with the node features. Neo-GNN (Yun et al., 2021) and BUDDY (Chamberlain et al., 2023) use heuristic functions to model high-order common neighbor information. NCNC (Wang et al., 2024) directly concatenates the weighted sum of node representations of common neighbors of $u$ and $v$ with the Hadamard product of MPNN representations of $u$ and $v$.
Graph Learning with Vision. Recent research has explored the potential of vision in graph learning. Das et al. (2023) find that the vision-language models have competitive node classification abilities. Wei et al. (2024) shows that visual perception benefits language-based graph reasoning tasks. However, no existing work realizes the potential of vision in enhancing MPNN for link prediction, which is the focus of this paper.
@@ 3 Preliminaries
Notations. A graph $G=(V,E)$ comprises of a set $V$ of $n$ nodes and a set $E$ of $e$ links. We denote the adjacency matrix of $G$ by $A\in\mathbb{R}^{n\times n}$, where $A_{(u,v)}>0$ if and only if the edge $(u,v)$ exists (i.e., $(u,v)\in E$). We define $N(v):=\{v\in V|A_{uv}>0\}$ as the set of neighbors of node $v$, and $N_{k}(v)$ as the set of $k$-hop neighbors of node $v$. In other words, a node $u\in N_{k}(v)$ if and only if $u$ and $v$ are reachable within $k$ hops. The node feature matrix $\mathbf{X_{G}}\in\mathbb{R}^{n\times F}$ contains the node features in $G$, where the $v$-th row $\mathbf{x_{v}}$ corresponds to the feature of node $v$. We use $S^{k}_{uv}=(V_{uv}^{k},E_{uv}^{k})$ to denote the $k$-hop subgraph enclosing the link $(u,v)$, where $V_{uv}^{k}$ is the union of $k$-hop neighbors of $u$ and $v$, and $E_{uv}^{k}$ is the union of links that can be reached by a $k$-hop walk originating at $u$ or $v$. Moreover, $S^{k}_{v}$ is the $k$-hop subgraph starting from node $v$.
MPNNs for Link Prediction. The MPNN is a commonly used model for link prediction. In MPNN, message passing is employed to iteratively update node representations based on information exchanged between neighboring nodes. Mathematically, this mechanism can be written as
@@
where $\bm{h}_{v}^{t}$ is the representation of node $v$ after $t$ layers, $U^{t}$, $M^{t}$ and $\mathbf{AGG}$ are the update, message-passing, and aggregation functions of layer $t$, respectively. $\mathbf{Y_{G}}\in\mathbb{R}^{n\times F^{\prime}}$ is the matrix of final node representations by MPNN for $G$, whose the $v$th row $\mathbf{y}_{v}$ is the final representation of $v$. Given $\mathbf{Y_{G}}$, link probabilities can be computed as $p(u,v)=R(\mathbf{y}_{u},\mathbf{y}_{v})$, where $R$ is a learnable readout function.
@@ 4 Methodology
In this section, we present the proposed GVN framework. An illustration of GVN is in Figure 1. In Sections 4.1 to 4.3, we will delve into the detailed designs inside, along with addressing the research questions raised in the introduction. In Section 4.4, we propose a computationally efficient variant of GVN for large-scale graphs.
@@ 4.1 Scope-Decoupled Subgraph Visualization
Link-Centered Subgraph Visualization. Beginning with RQ1 (i.e., For the link prediction task, what is the effective approach to be aware of graph structures from vision?), we transform graph structures into visual representations by rendering them as images during pre-processing. As large-scale graphs often have thousands of nodes and edges, we utilize $k$-hop subgraph sampling, a common practice in graph learning (Hamilton et al., 2017). This method concentrates on visual awareness via the most relevant local $k$-hop subgraph $S^{k}_{uv}$ around the central link $(u,v)$, while excluding distant structures that may introduce noise. In line with the locality principle [^1] Locality Principle: Correlated information is contained within the near neighbors (Bronstein et al., 2017; Wu et al., 2020). and effective practices of many MPNNs with $k\leq 3$ (Kipf & Welling, 2017; Hamilton et al., 2017; Xu et al., 2019; Velickovic et al., 2018; Chen et al., 2022), we limit $k$, which defines the visualized scope, to a maximum of 3.
To be specific, utilizing existing automated graph visualization tools (Details in Appendix F) such as Graphviz (Gansner & North, 2000), Matplotlib (Tosi, 2009), and Igraph (Gabor Csardi, 2006), a given query link $(u,v)$ and its surrounding $k$-hop subgraph $S^{k}_{uv}$ are mapped one-to-one into an image $I^{k}_{uv}$ by a graph visualizer $\mathrm{GV}$. Formally, this is expressed as $I^{k}_{uv}=\mathrm{GV}(S^{k}_{uv},u,v)$.
@@
1) Keeping Style Consistency. It is crucial that all visualized subgraphs maintain consistency in the graph visualizer $\mathrm{GV}$, such as uniformly using Graphviz with fixed configurations (e.g., graph layout algorithm, node colors, and shapes). This prevents inconsistent image styles from introducing additional learning difficulties and unstable training.
@@
3) Ablating Node Labels. All the node labels are omitted, enabling the model to focus on the graph structure, which benefits generalizability.
Figure 2: Illustrations of visual graph images with different $k$’s.
Decoupled Vision Scope. While the perception scope expands during message passing in MPNN, we fix the visual perception scope $k$ during visualization. This ensures that the scope in vision can be decoupled from which in message passing and offers two key advantages. First, both theoretical analysis and empirical evidence (Zeng et al., 2021) show that a decoupled scope independent from message-passing can improve expressivity of MPNN by addressing the over-smoothing issue (Li et al., 2018; Kenta Oono, 2020; Shen et al., 2024), and enhances the computational scalability of the models by alleviating the neighbor explosion problem (Chiang et al., 2019; Zeng et al., 2019). Second, as shown in Figure 2, an excessively large visual perception scope can lead to overcrowding of nodes/edges on a fixed-size canvas, resulting in a cluttered appearance that complicates graph interpretation. By using a small $k$, we maintain image clarity independent of message passing.
@@ 4.2 Adaptive Extraction of Visual Structural Features
With the same perceptive scope, VSFs can discriminate links involving isomorphic nodes, whereas MPNNs or the 1-dimensional Weisfeiler-Lehman test cannot.
@@
Fine-grained Substructure Awareness. The ability to capture substructures (motifs) is another perspective of MPNN expressive power and plays a crucial role in areas like biology, molecular, and social networks (Chen et al., 2020; Kanatsoulis & Ribeiro, 2024; Yan et al., 2024). However, this can be challenging for MPNNs (Chen et al., 2020; Arvind et al., 2020). We demonstrate in the following that VSFs are helpful for MPNNs in capturing substructures. Specifically, we follow the experimental setup in (Chen et al., 2020). This involves generating two types of graph datasets, Erdős-Rényi graphs and random regular graphs, and compute the counts of two substructures including triangles and 3-stars as ground-truth labels. We use MPNN followed by a 3-layer MLP decoder as the architecture and the VSFs are concatenated after the MPNN-produced representations. The evaluation metric is the normalized mean squared error (i.e., ratio of the mean squared error over the variance of ground-truth labels). Table 1 shows the best and median (third-best) performance over five runs with different random seeds. As can be seen, incorporating VSFs significantly improves the substructure capturing abilities of MPNNs. More details of this experiment are provided in Appendix E.
@@
Information-rich and Adaptive SF Extraction. While the design of existing structural features rely on heuristics, VSFs encode the whole subgraph, making them more information-rich. Figure 4 shows the reproduction ratios [^3] The reproduction ratio is defined as the percentage (%) of SFs successfully predicted. of various SFs based on VSFs on ogbl-ddi, by feeding the VSFs extracted from the pretrained ResNet50 into a trainable 3-layer MLP. As can be seen, VSFs effectively capture most of the information from a wide array of structural features, underscoring their comprehensive nature as a versatile structural feature pool.
@@
The reproduction ratios of various SFs reflect the relative distribution of information within the VSFs. Table 2 compares the reproduction ratios of SFs between VSFs obtained from a pre-trained ResNet50 and those after finetuning. The latter is derived from a ResNet50 encoder fine-tuned for link prediction with a 2-layer MLP on the ogbl-ddi and PubMed datasets. Note that ogbl-ddi is a very dense graph, while PubMed is much sparser. As shown in Table 2, after finetuning on ogbl-ddi, the reproduction ratio of path-based SFs (SPD, DRNL, and DE) consistently decreases. This aligns with the under-performance observed in path-based link prediction models like SEAL (Zhang & Chen, 2018) and NBFNet (Zhu et al., 2021) on the ogbl-ddi dataset. This is because in the ogbl-ddi dataset, most node pairs are reachable within two hops, making path-based structural features like SPD less informative. On the other hand, in the sparser PubMed dataset where path information is more valuable, the reproduction ratios of common neighbor-based similarity functions (CN, RA, AA) consistently increase after finetuning. This trend is demonstrated by models such as NeoGNN, Buddy, and NCCN (Yun et al., 2021; Chamberlain et al., 2023; Wang et al., 2024), which consistently benefit from common neighbor similarity functions across various datasets. These observations highlight the ability of VSFs to adapt the internal information types to meet the specific demands of different scenarios through finetuning.
@@
Table 2: Reproduction ratios of SFs in VSFs with and without finetuning on the ogbl-ddi and PubMed datasets.
@@
Through finetuning, VSFs can adaptively adjust their emphasis on different types of structural information, thus providing scenario adaptability.
VSFs are not permutation-equivariant, yet they remain effective with MPNNs, as will be shown in Section 5.1. This is potentially attributed to two reasons: 1) Learnable VSFs capture specific information like structural patterns or motifs (see Remark 4.2). 2) Visualization introduces permutation-based augmentations (e.g., varying layouts), making the decoder less sensitive to node order.